OpenStudy (anonymous):

Together, teammates Pedro and Ricky got 2685 base hits last season. Pedro had 279 more hits than Ricky. How many hits did each player have? using Solving Systems of Linear Equations Using Substitution

OpenStudy (anonymous):

2865-279=2586

OpenStudy (anonymous):

2586/2 = 1293

OpenStudy (anonymous):

say pedro is one half and ricky the other

OpenStudy (anonymous):

put that 279 back in for pedro so 1293+279=

OpenStudy (anonymous):

answer: 1572

OpenStudy (anonymous):

answer: 1572 is right

OpenStudy (arabpride):

LOL #copypaste